# Dynamical Formation of Apparent Horizons due to Boundary Effect in Vacuum Einstein Gravity

### v2

@@ 3.1. Main Ideas and Novelty
The guiding principle is therefore to drive the difference $\mbox{tr}\chi-\mbox{tr}\underline{\chi}$ to become large and positive while simultaneously suppressing the absolute term $|\mbox{tr}\chi+\mbox{tr}\underline{\chi}|$ along the incoming null direction.
This strategy is realized by prescribing a large incoming shear $\underline{\hat{\chi}}$ on the initial null hypersurface $\underline{H}_{0}$ and an appropriately scaled profile for $\mbox{tr}\chi$ and $\mbox{tr}\underline{\chi}$ on the distant outgoing hypersurface $u=u_{\infty}$ consistent with the hierarchy of the data. The evolution of $\mbox{tr}\underline{\chi}$ along the incoming direction is then governed by its transport equation, which contains both quadratic and indefinite–sign nonlinearities. The main difficulty arises in ensuring that the positive enforcing term
$a^{-1}\int_{u_{\infty}}^{-a}|u^{\prime}|\,|\underline{\hat{\chi}}|^{2}\,du^{\prime}$
dominates all lower–order contributions. To achieve this dominance, we impose a relative largeness condition
(48) $\Gamma^{0}_{\underline{\hat{\chi}},2}>100\,\Gamma^{0}_{2},$
where $\Gamma^{0}_{\underline{\hat{\chi}},2}$ and $\Gamma^{0}_{2}$ denote the scale–invariant initial data norms controlling $\underline{\hat{\chi}}$ and the remaining Ricci coefficients, respectively. Remarkably, this relative hierarchy is propagated uniformly throughout the entire semi-global development for sufficiently large parameter $a$ used in this study.
An additional constraint arises from the requirement that the radius–mean curvature inequality (9) is satisfied along evolution. The mechanism can be schematically summarized as follows. Along the incoming direction, we enforce $\mbox{tr}\underline{\chi}$ to grow in magnitude near u=au=-a, while $\mbox{tr}\chi$ remains uniformly positive. Owing to the monotonicity property implied by (45), the only viable means of maintaining positivity of $\mbox{tr}\chi$ is to restrict the $\underline{u}$–width of the slab to be sufficiently small. Geometrically, this ensures that the t=at=-a Cauchy slice bounded by $S_{-a,0}$ possesses a smaller radius than its nearby deformation bounded by $S_{-a,\epsilon}$ on $t=-a+\epsilon$, as required for the desired curvature inequality.
The methodology presented here represents a substantial conceptual departure from all previous works on trapped surface formation. The open set of characteristic data constructed herein accommodates a broader class of large data than those treated in the earlier frameworks of Christodoulou and Klainerman–Rodnianski. Moreover, the Cauchy data prescribed on the interior slice $\mathcal{M}_{1}$ is defined implicitly through a Riemannian construction in the spirit of Yau [38], involving the notion of prescribed Yau radius and mean curvature control. This interior data will be rendered explicit and developed further in the companion paper [28].
The analytical framework developed here constitutes a genuinely new approach to the dynamical formation of apparent horizons and trapped surfaces in vacuum general relativity. To the best of our knowledge, no previous result in the literature establishes the emergence of such geometric features through the control of the generalized mean curvature quantity $c=H-|\kappa|$ within a semi–global evolution. In particular, the present construction yields an open set of characteristic initial data exhibiting a substantially broader admissible class than those considered in the works of Christodoulou, Klainerman–Rodnianski, or Luk, thereby enlarging the known landscape of large–data configurations leading to horizon formation.
The corresponding Cauchy data prescribed on the interior slice $\mathcal{M}_{1}$ is, at this stage, defined only implicitly through a Riemannian geometric construction inspired by Yau [38]. Specifically, the data $(g,k)$ on $\mathcal{M}_{1}$ are required to satisfy the boundary mean curvature condition consistent with a prescribed Yau radius, ensuring the geometric compatibility between the interior region and the exterior double–null development. A concrete realization of this construction, including an explicit parametrization of $(g,k)$ and verification of the constraint equations, will be presented in the sequel [28].
A fundamental analytical difficulty arises in propagating the full hierarchy of weighted and scale-invariant data norms from the initial null and Cauchy hypersurfaces throughout the domain of semi-global existence. Once this propagation is established, the Riemannian argument of Yau [38], when applied to the interior slice $\mathcal{M}_{1}$, becomes decisive: the boundary mean curvature condition then determines, in a geometric and quantitative sense, whether an apparent horizon must or must not exist. It is noteworthy that the scaling framework introduced by An [3], after a slight modification consistent with the present gauge choice, integrates seamlessly into this construction and provides the correct normalization for the hierarchy of Ricci coefficients and curvature fluxes.
Another distinctive aspect of the present analysis is that elliptic estimates are not required. By formulating the entire problem at a sufficiently high level of Sobolev regularity within the hyperbolic (double-null) setting, we circumvent any dependence on auxiliary elliptic estimates and maintain full control through transport-type and energy-type arguments (this is well known by now in the context of the characteristic initial value problem). Finally, the structure developed here aligns naturally with the local constructions of Luk [24], suggesting that one can generate more general families of initial data on finite domains satisfying Yau’s boundary condition in a dynamically consistent fashion.
4. Setup
@@ 4.1. Construction of the double null gauge
Denote by $L\mathcal{M}$ the frame bundle of $\mathcal{M}$. We construct a double null gauge, meaning a smooth section of this bundle such that, through it, each point $p\in\mathcal{M}$ maps to a renormalized frame $(e_{1},e_{2},e_{3},e_{4})\in L\mathcal{M}$ with $g(e_{3},e_{4})=-2$, $g(e_{A},e_{B})=\epsilon_{AB}$ and $g(e_{3},e_{A})=g(e_{4},e_{A})=0$.
We begin with two null hypersurfaces $H_{u_{\infty}},\underline{H}_{0}$ and their intersection $S_{u_{\infty},0}$, a topological 2-sphere. For any point $q$ on this 22-sphere, the tangent space $T_{q}S_{u_{\infty},u}$ is 2-dimensional and admits a 22-dimensional orthogonal complement $T_{q}^{\perp}{S_{u_{\infty},0}}$, on which we can find two future-directed null vectors $L^{\prime}_{q}$ and $\underline{L}^{\prime}_{q}$, normalized so that
@@ 4.2. Choice of coordinates and expression of the metric
It is a requirement that $b^{A}\equiv 0$ on $\underline{H}_{0}$. The section that maps $p\in\mathcal{M}\mapsto\left({\left.\kern-1.2pt\theta^{1}\right|_{p}},{\left.\kern-1.2pt\theta^{2}\right|_{p}}{\left.\kern-1.2pte_{3}\right|_{p}},{\left.\kern-1.2pte_{4}\right|_{p}}\right)$ is the double null gauge we wanted to construct.
@@ 4.3. The vacuum Einstein equations expressed in the double null gauge
In this section we are going to express the Einstein-Yang-Mills system in the double null gauge given above. We begin by decomposing curvature components and Ricci coefficients with respect to the frame $(e_{1},e_{2},e_{3},e_{4})$. Let $A,B$ take values in $\begin{Bmatrix}1,2\end{Bmatrix}$. We define the following (Weyl) null curvature components:
For the Ricci coefficients, we decompose as follows:
$\zeta_{A}:=\frac{1}{2}g(D_{A}e_{4},\hskip 1.42262pte_{3}).$
@@
Here $\not{\epsilon}$ is the volume form associated with the metric $\gamma$. Moreover, by $\phi^{1}\cdot\phi^{2}$ we mean an arbitrary contraction of the tensor product of $\phi^{1}$ and $\phi^{2}$ with respect to the metric $\gamma$. We also define suitable trace, divergence and curl operators. For totally symmetric tensors $\phi$, we define these operators as follows:
$(\mbox{div }\phi)_{A_{1}\dots A_{r}}:=\nabla^{B}\phi_{BA_{1}\dots A_{r}},$
$(\mbox{curl }\phi)_{A_{1}\dots A_{r}}:=\not{\epsilon}^{BC}\nabla_{B}\phi_{CA_{1}\dots A_{r}},$
$(\mbox{tr}\phi)_{A_{1}\dots A_{r-1}}:=(\gamma^{-1})^{BC}\phi_{BCA_{1}\dots A_{r-1}}.$
@@
(65) $K-\frac{1}{2}\hat{\chi}\cdot\hat{\underline{\chi}}+\frac{1}{4}\mbox{tr}\chi tr\underline{\chi}=-\rho,$
(66) $\nabla_{4}\zeta=2\nabla\omega+\chi\cdot(\underline{\eta}-\zeta)+2\omega(\zeta+\underline{\eta})-\beta,$
(67) $\nabla_{3}\zeta=-2\nabla\omega-\underline{\chi}\cdot(\zeta+\eta)+2\underline{\omega}(\zeta-\eta)-\underline{\beta}$
@@ 4.4. Integration
For a spacetime region $\mathcal{D}_{u,\underline{u}}:=\begin{Bmatrix}\left(u^{\prime},\underline{u}^{\prime},\theta^{1},\theta^{2}\right)\hskip 1.42262pt\mid\hskip 1.42262ptu_{\infty}\leq u^{\prime}\leq u,0\leq\underline{u}^{\prime}\leq\underline{u}\end{Bmatrix}$, we define the spacetime integral
(81) $\int_{\mathcal{D}_{u,\underline{u}}}\phi:=\sum_{U}\int_{u_{\infty}}^{u}\int_{0}^{\underline{u}}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\phi\cdot p_{U}\cdot\Omega^{2}\cdot\sqrt{-\text{det}g}\hskip 1.42262pt\text{d}\theta^{1}\text{d}\theta^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\hskip 1.42262pt\text{d}u^{\prime}.$

### v4

@@ 2.5. Comparison with Previous Studies and Novelty
The guiding principle is therefore to drive the difference $\mbox{tr}\chi-\mbox{tr}\underline{\chi}$ to become large and positive while simultaneously suppressing the absolute term $|\mbox{tr}\chi+\mbox{tr}\underline{\chi}|$ along the incoming null direction. This is essential not to simply control the generalized mean curvature $c=H-|\kappa|$, but also to control the radius of the domain under consideration (and gluing) as seen in the section 2.4 of introduction.
The construction is implemented by prescribing the incoming shear $\underline{\hat{\chi}}$ on the initial incoming null hypersurface $\underline{H}_{0}$ together with a normalized dispersive profile for the null expansions $\mbox{tr}\chi$ and $\mbox{tr}\underline{\chi}$ on the distant outgoing hypersurface $H_{u_{\infty}}$. The data are arranged so as to satisfy a precise scale–invariant hierarchy which departs substantially from the Minkowskian regime while retaining dispersive decay in $|u|$. The parameter $a\gg 1$ measures the amplitude of the deviation from flat data, whereas inverse powers of $|u|$ encode null dispersion along the foliation.
More precisely, the Ricci coefficients are assumed to obey the asymptotic hierarchy
$\bigg|\mbox{tr}\chi-\frac{2}{|u|}\bigg|=O\!\left(a^{-\frac{1}{2}}|u|^{-1}\right),\qquad|\hat{\chi}|=O\!\left(a^{-\frac{1}{2}}|u|^{-1}\right),\qquad\bigg|\mbox{tr}\underline{\chi}+\frac{2}{|u|}\bigg|=O\!\left(|u|^{-2}\right),\qquad|\underline{\hat{\chi}}|=O\!\left(a^{\frac{1}{2}}|u|^{-2}\right),$
$|\omega|=O\!\left(a^{\frac{1}{2}}|u|^{-2}\right),\qquad|\underline{\omega}|=O\!\left(a^{\frac{1}{2}}|u|^{-3}\right),\qquad|\eta|+|\underline{\eta}|=O\!\left(a^{\frac{1}{2}}|u|^{-2}\right).$
This scaling is consistent with the null structure equations and is chosen so that the dominant large component is the incoming shear, while all remaining Ricci coefficients remain perturbative relative to their Minkowskian leading orders.
A principal analytic difficulty is that the null transport equations for the Ricci coefficients contain Weyl curvature components as source terms. These curvature components must satisfy decay estimates in $|u|$ compatible with the peeling behavior (cf. [26]) and, simultaneously, an amplitude hierarchy compatible with the above Ricci coefficient scaling. In particular, the curvature components are arranged to satisfy
$|\alpha|=O\!\left(a^{-\frac{1}{2}}|u|^{-1}\right),|\beta|=O\!\left(|u|^{-2}\right),|\underline{\beta}|=O\!\left(a|u|^{-4}\right),|\rho|+|\sigma|=O\!\left(a^{\frac{1}{2}}|u|^{-3}\right),|\underline{\alpha}|=O\!\left(a^{\frac{3}{2}}|u|^{-5}\right).$
The hierarchy is closed in the sense that, under the null structure and Bianchi equations, these weights are stable under propagation in the semi–global region of existence up to controlled losses. Ensuring this compatibility and propagation of scale is a central structural requirement in the argument.
Remark 7.
Note how the largeness of $a$ manifests. In simpler geometries such as an asymptotically Schwarchild case, $\rho=O(m|u|^{-3})$ and so $a^{\frac{1}{2}}$ in our context roughly behaves like mass-in fact the sphere $S_{-a,0}$ has Hawking mass $\approx a^{1/2}$. Thus, for $a\gg 1$, the data manifestly falls under the category of the large data or in the moduli space, $a^{\frac{1}{2}}$ essentially measures a notion of ‘distance’ from the Minkowski space. Contrast this scaling with the previous works [5, 4, 12], where $a$ appears instead of $a^{\frac{1}{2}}$ and the scaling of the radiation fields are different.
It is instructive to compare the incoming gravitational energy in the short–pulse framework of Christodoulou and its later refinements (see [19, 5, 4]) with the present construction. In the double–null formalism, the incoming gravitational energy flux through an outgoing null hypersurface $H_{u_{\infty}}$ over the slab $\underline{u}\in[0,1]$ is measured by
$E_{\infty}:=\int_{0}^{1}\!\int_{S_{u_{\infty},\underline{u}}}|\hat{\chi}|^{2}\,d\mu_{\gamma}\,d\underline{u},\qquad d\mu_{\gamma}=\sqrt{\det(\gamma_{AB})}\,.$
In the short–pulse regime of [5, 4], one prescribes
$|\hat{\chi}(u_{\infty},\underline{u})|\sim a^{1/2}|u_{\infty}|^{-1},\qquad\gamma(u_{\infty},\underline{u})\sim|u_{\infty}|^{2}\gamma_{0},$
so that $|S_{u_{\infty},\underline{u}}|\sim|u_{\infty}|^{2}$ and therefore
$E_{\infty}^{\mathrm{short\text{-}pulse}}\;\approx\;\int_{0}^{1}\!\int_{S_{u_{\infty},\underline{u}}}a\,|u_{\infty}|^{-2}\,d\mu_{\gamma}\;\sim\;a.$
Thus the incoming gravitational energy from past null infinity is large.
In contrast, in the present hierarchy one imposes the dispersive scaling
$|\hat{\chi}(u_{\infty},\underline{u})|\sim a^{-1/2}|u_{\infty}|^{-1},$
with the same area scale for $S_{u_{\infty},\underline{u}}$. Consequently,
$E_{\infty}^{\mathrm{present}}\;\approx\;\int_{0}^{1}\!\int_{S_{u_{\infty},\underline{u}}}a^{-1}|u_{\infty}|^{-2}\,d\mu_{\gamma}\;\sim\;a^{-1},$
which is small. The present construction therefore operates in a regime of weak incoming radiation from null infinity, in sharp contrast with the short–pulse mechanism.
We next examine the Hawking mass of the boundary sphere $S_{-a,0}\subset\underline{H}_{0}$. Recall that
$m_{H}(S)=\frac{r}{2}\left(1+\frac{1}{16\pi}\int_{S}\mbox{tr}\chi\,\mbox{tr}\underline{\chi}\,d\mu_{\gamma}\right),\qquad r=\Big(\frac{|S|}{4\pi}\Big)^{1/2}.$
On $S_{-a,0}$ we have $r\sim a$ and, by the Gauss equation,
$K=-\rho-\frac{1}{4}\mbox{tr}\chi\,\mbox{tr}\underline{\chi}+\frac{1}{2}\hat{\chi}\cdot\underline{\hat{\chi}}$.
Using $K\sim a^{-2}$ and $|\hat{\chi}\cdot\underline{\hat{\chi}}|\ll a^{-2}$ under the assumed hierarchy, one obtains
$\int_{S_{-a,0}}\mbox{tr}\chi\,\mbox{tr}\underline{\chi}\,d\mu_{\gamma}=-4\int_{S_{-a,0}}\rho\,d\mu_{\gamma}+O(1).$
With $|\rho|\sim a^{1/2}|u|^{-3}\sim a^{-5/2}$ at u=au=-a and $|S_{-a,0}|\sim a^{2}$, it follows that
$\int_{S_{-a,0}}\rho\,d\mu_{\gamma}\sim a^{-1/2},\qquad m_{H}(S_{-a,0})\sim a^{1/2}\gg 1.$
Thus the Hawking mass of the initial spheres foliating $\underline{H}_{0}$ is large despite the small incoming radiation from $H_{u_{\infty}}$.
This should be contrasted with the characteristic setup in Christodoulou’s short–pulse framework [19], where the incoming hypersurface $\underline{H}_{0}$ is taken to be exactly Minkowskian and all large effects arise from concentrated incoming radiation. The mechanism realized here is different: it is not a collapse driven by a high–energy short pulse, but a boundary–driven large–scale configuration in which substantial mass is distributed over an isotropically large domain while outgoing radiation remains mild.
This viewpoint is consistent with the boundary–driven geometric mechanisms established by Schoen and Yau [38] in the Riemannian setting, where quantitative lower bounds on boundary mean curvature impose rigid constraints on the admissible interior geometry and topology. In that context, sufficiently strong boundary convexity forces global interior consequences independent of any concentration of bulk energy. An analogous phenomenon was later identified by Yau in the Lorentzian setting, where appropriate boundary geometric conditions guarantee the existence of marginally outer trapped surfaces at the level of initial data, without requiring positive matter density assumptions.
The present analysis gives a genuinely dynamical realization of this boundary–effect principle within the Einstein vacuum equations. The mechanism developed here does not rely on short–pulse–type concentration of incoming gravitational radiation. Instead, the decisive inputs are the persistence of strong boundary mean curvature along the outgoing null hypersurface and the presence of large total mass distributed over an isotropically large spatial domain. Under these conditions, the null evolution amplifies the boundary geometry in a controlled manner and leads to the formation of a marginally outer trapped surface in the interior of the spacetime development.
In particular, this establishes that black hole formation is not restricted to collapse scenarios driven by highly concentrated energy flux. It also occurs in a complementary large–scale regime characterized by dispersed mass and mild radiation, thereby identifying a distinct and robust geometric pathway to trapped surface formation.
A key structural feature in our argument is the isotropic largeness of the strong–field region, quantified through a lower bound for the Schoen–Yau (HH–) radius. What is essential is not largeness of total mass alone, but largeness of the domain on which curvature and mean–convexity are simultaneously controlled in an essentially direction–independent manner.
This distinction is clarified by the localized gluing construction of Carlotto–Schoen [14]. Their theorem produces asymptotically flat, scalar–flat (time–symmetric) vacuum initial data sets with arbitrarily large ADM mass whose geometry agrees with a large–mass Schwarzschild end inside a prescribed cone, while remaining exactly Euclidean outside a slightly larger cone. In particular, the gravitational field can be made strong yet highly anisotropic, being confined to a narrow angular sector and completely shielded elsewhere.
Such examples show that large ADM mass by itself does not enforce any uniform, isotropic geometric control on large coordinate balls or quasi-round domains. In particular, one cannot deduce from mass alone the presence of a large mean–convex barrier region or a domain with large Schoen–Yau radius to which boundary–driven minimal or trapped surface arguments apply. The obstruction is geometric: curvature concentration that is strongly directional can be separated from large portions of the manifold by exactly flat regions.
This behavior stands in sharp contrast with the regime considered here. Our hypotheses impose quantitative mean–curvature and radius control on an isotropically large domain, ensuring that the dominant curvature and energy flux are not confined to a thin sector but are distributed over a region with uniform geometric thickness. It is precisely this combination — large scale together with strong positive boundary generalized mean curvature-that allows one to convert curvature concentration into a marginally outer trapped surface through evolution. The result, therefore, isolates a mechanism for MOTS formation that depends on global geometric size and boundary convexity, rather than on total mass alone or quasi-local energy concentration.
Figure 3. An equivalent alternative configuration yielding MOTS in an evolutionary manner. One may technically not need the characteristic evolution that is performed in this work. Recall that the data on middle part $\mathcal{M}_{2}$ is induced by the characteristic development. However, since we already know the type of data induced on $\mathcal{M}_{2}$, one can in principle start with this data on $\mathcal{M}_{2}$, the interior data on $\mathcal{M}_{1}$ and Kerr exterior on the outside $\mathcal{M}_{3}$ and initiate the time evolution. By uniqueness of the solution of the Cauchy problem of the Einstein’s equations one would have MOTS in the interior of isotropically large domain $\mathcal{M}^{{}^{\prime}}_{1}$.
2.6. Acknowledgements
This work is supported by the Beijing Institute of Mathematical Sciences and Applications of the Yau Mathematical Science Center at Tsinghua University and Beijing Municipal Govt.
3. Part 1: Solution of the Characteristic Part
This is an important aspect of this study. But the study of characteristic initial value problems and proving a semi-global development with appropriate smooth initial data is well-established. We will follow the signature of decay rate technology developed by An [5] that is designed to handle large data regime in a systematic way.
@@ 3.1. Construction of the double null gauge
Denote by $L\mathcal{M}$ the frame bundle of $\mathcal{M}$. We construct a double null gauge, meaning a smooth section of this bundle such that, through it, each point $p\in\mathcal{M}$ maps to a renormalized frame $(e_{1},e_{2},e_{3},e_{4})\in L\mathcal{M}$ with $g(e_{3},e_{4})=-2$, $g(e_{A},e_{B})=\delta_{AB}$ and $g(e_{3},e_{A})=g(e_{4},e_{A})=0$.
We begin with two null hypersurfaces $H_{u_{\infty}},\underline{H}_{0}$ and their intersection $S_{u_{\infty},0}$, a topological 2-sphere. For any point $q$ on this 22-sphere, the tangent space $T_{q}S_{u_{\infty},u}$ is 2-dimensional and admits a 22-dimensional orthogonal complement $T_{q}^{\text{Per}p}{S_{u_{\infty},0}}$, on which we can find two future-directed null vectors $L^{\prime}_{q}$ and $\underline{L}^{\prime}_{q}$, normalized so that
@@ 3.2. Choice of coordinates and expression of the metric
The section that maps $p\in\mathcal{M}\mapsto\left({\left.\kern-1.2pt\theta^{1}\right|_{p}},{\left.\kern-1.2pt\theta^{2}\right|_{p}}{\left.\kern-1.2pte_{3}\right|_{p}},{\left.\kern-1.2pte_{4}\right|_{p}}\right)$ is the double null gauge we wanted to construct.
@@ 3.3. The vacuum Einstein equations expressed in the double null gauge
In this section we are going to express the Einstein equations in the double null gauge given above.
@@
Here $\not{\epsilon}$ is the volume form associated with the metric $\gamma$. Moreover, by $\phi^{1}\cdot\phi^{2}$ we mean an arbitrary contraction of the tensor product of $\phi^{1}$ and $\phi^{2}$ with respect to the metric $\gamma$. We also define suitable trace, divergence, and curl operators. For totally symmetric tensors $\phi$, we define these operators as follows:
$(\mbox{div }\phi)_{A_{1}\dots A_{r}}:=\nabla^{B}\phi_{BA_{1}\dots A_{r}},(\mbox{curl }\phi)_{A_{1}\dots A_{r}}:=\not{\epsilon}^{BC}\nabla_{B}\phi_{CA_{1}\dots A_{r}},(\mbox{tr}\phi)_{A_{1}\dots A_{r-1}}:=(\gamma^{-1})^{BC}\phi_{BCA_{1}\dots A_{r-1}}.$
@@
(81) $K-\frac{1}{2}\hat{\chi}\cdot\hat{\underline{\chi}}+\frac{1}{4}\mbox{tr}\chi tr\underline{\chi}=-\rho.$
@@ 3.4. Integration
For a spacetime region $D_{u,\underline{u}}:=\begin{Bmatrix}\left(u^{\prime},\underline{u}^{\prime},\theta^{1},\theta^{2}\right)\hskip 1.42262pt\mid\hskip 1.42262ptu_{\infty}\leq u^{\prime}\leq u,0\leq\underline{u}^{\prime}\leq\underline{u}\end{Bmatrix}$, we define the spacetime integral
(95) $\int_{D_{u,\underline{u}}}\phi:=\sum_{U}\int_{u_{\infty}}^{u}\int_{0}^{\underline{u}}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\phi\cdot p_{U}\cdot\Omega^{2}\cdot\sqrt{-\text{det}g}\hskip 1.42262pt\text{d}\theta^{1}\text{d}\theta^{2}\hskip 1.42262pt\text{d}\underline{u}^{\prime}\hskip 1.42262pt\text{d}u^{\prime}.$